OpenStudy (anonymous):

Thermal physics! how do you rearrange the formula (q=mc ΔT) to the temp when you have everything except for the second temp, for the ΔT?

OpenStudy (anonymous):

T2-T1 =q/mc or T2=T1 + q/mc